OpenStudy (anonymous):

ABSOLUTE VALUE Solve. for x. | -8x - 9 | = 8 Step by step please?

OpenStudy (anonymous):

When you have absolute value, whatever is in the | | it will ALWAYS be positive .

OpenStudy (anonymous):

o.o why'd you leave...?

OpenStudy (anonymous):

It's ok... I just wanted clarification on whether or not I did this correctly? -8x-9=8 or -8x-9=-8 -8x=17 or -8x=1 x=-17/8 or x=-1/8

OpenStudy (anonymous):

-8x-9=8 and -8x-9=-8 solve both one by one. you are doing well now solve other also.

OpenStudy (anonymous):

I don't see what's left to be solved unless you're talking about the fractions?

OpenStudy (anonymous):

Ooooh ok lol

OpenStudy (anonymous):

I'll brb

OpenStudy (anonymous):

sorry you have solved both and they are correct. Don't solve sde by side .solve one by one.

OpenStudy (anonymous):

when x=-17/8 -(-17/8)*8-9=17-9=8 so absolute value=8 similarly whenx=-1/8 -(-1/8)*8-9=1-9=-8 again absolute value is 8

OpenStudy (anonymous):

is everything clear?

OpenStudy (anonymous):

Ooooh I see... thank you very much.

OpenStudy (anonymous):

you are welcome.